OpenStudy (anonymous):

integrate ln sqrt (x^2-1)

OpenStudy (saifoo.khan):

Ops, it's ln.

OpenStudy (anonymous):

oh and i suppose if i take the derivative of that i will get \[\ln(\sqrt{x^2-1})\] fat chance!

OpenStudy (saifoo.khan):

so the integration for ln = 1/x im not really sure. i have to check my register for it. :/ @LagrangeSon678

OpenStudy (anonymous):

its gonna be alittle harder than that

OpenStudy (anonymous):

it looks like intergration by parts. Is that the best way to go with this problem?

OpenStudy (anonymous):

thats seems like a good route, alhtough i'll leave the work up to you

OpenStudy (saifoo.khan):

i got it. :D

OpenStudy (saifoo.khan):

wait. i got the derivatives of ln. -_-

OpenStudy (saifoo.khan):

@LagrangeSon678 , please continue.

OpenStudy (anonymous):

\[\ln(\sqrt{x^2-1})=\frac{1}{2}\ln(x-1)+\frac{1}{2}\ln(x+1)\] is a good start, also good is to have memorized \[\int \ln(x)dx=x\ln(x)-x\] otherwise it is groundhog day all over again

OpenStudy (anonymous):

don't forget to "simplfy" using properties of the logs to start. makes life much easier. the rest you can pretty much do in your head so long as you can remember what the anti derivative of the log is. memorize it, because math teachers love to put in on tests

OpenStudy (anonymous):

thanks. for the help